OpenStudy (anonymous):

prove sinx over 1-cosx= 1+cosx over sinx

OpenStudy (lalaly):

\[\frac{sinx}{1-cosx} \times \frac{1+cosx}{1+cosx}=\frac{sinx(1+cosx)}{1-\cos^2x}\]\[=\frac{sinx(1+cosx)}{\sin^2x}=\frac{1+cosx}{sinx}\]

OpenStudy (anonymous):

sin^3x+cos^3x over sinx+cosx = 1-sinx cosx